Calculate the mole fraction of each component of a solution containing 65 g of ethanol

(C2H6O) in 350 g of water.


1
Expert's answer
2021-02-23T04:36:19-0500

n1= 65/46 =1,41304348 mol of C2H6O.

n2= 350/18 = 19,4444444 mol of H2O.

w%(mol fraction) = n1x100%/n1+n2= 1,41304348 x 100%/1,41304348 + 19.4444444 = 6,77475392% of C2H6O.

93,2252461 % of H2O.